Hi, I had just learned about probability, premutaions and combinations, but I am a bit confused with everything. I find it difficult to understand when to use which and how to use it. So I need help in a question I was thinking of...
The question: Suppose I am one of 75 people who are contestents to win a free traveling ticket. There are 25 winning places. What is the probability of me being chosen as a winner.
I don't think we should use a premutation, but I can;t about anything else about how to solve this question.
Thanks in advance for your help.

Answer by rothauserc(4718)

there are 75 contestants and 25 winning places, therefore in the first drawing you have 25/75 = 1/3, which means a one in three chance on the first draw
assuming you keep having drawings until all winning places have been determined
(1/3)+(24/74)+(23/73)+(22/72)+(21/71)+(20/70)+(19/69)+(18/68)+(17/67)+(16/66)+(15/65)+(14/64)+(13/63)+(12/62)+(11/61)+(10/60)+(9/59)+(8/59)+(7/58)+(6/57)+(5/57)+(4/56)+(3/55)+(2/54)+(1/53) = your probability of winning one of the winning places